# Illuminating Darkness: Enhancing Real-world Low-light Scenes with Smartphone Images

### v1

@@ 4.1 Comparison between SLLIE methods
We evaluated state-of-the-art (SOTA) methods to summarize their performance on real SLLIE. This study focused on Retinex-based models, including RetinexNet [49], Kind [61], Kind+ [62], and DeepUPE [47], as well as transformer-based SLLIE methods like Retinexformer [4], IAT [7], and SNRNet [52]. We also assessed the capabilities of SOTA restoration models, Uformer [48], and MIRNet [57], to evaluate their applicability in real-world SLLIE scenarios. Please note that we included methods that leverage Retinex theory or transformer architectures for fair comparison. We verified the reported results from their respective papers through our implementations using the recommended datasets before training them with the proposed LSD. To further generalize our findings, we trained all models on a combined dataset comprising both DLI and NLI images, ensuring a robust evaluation across diverse real-world conditions.
Quantitative Evaluation. Tab. 2 presents the performance of our SLLIE methods on the LSD dataset, covering both indoor and outdoor scenes captured under low-light and extreme low-light conditions. Our evaluation demonstrates that the proposed TFFormer consistently outperforms SOTA methods under all conditions, including denoised (DLI) and noisy (NLI) scenarios. In comparison to the most substantial baseline, RetinexFormer [4], TFFormer delivers improvements of 1.6 dB in PSNR, 0.0635 in SSIM, and a reduction of 0.021 in LPIPS on the DLI subset, as well as gains of 2.13 dB in PSNR, 0.0745 in SSIM, and a reduction of 0.0509 in LPIPS on the NLI subset. On average, TFFormer achieves enhancements of 1.865 dB in PSNR, 0.138 in SSIM, and a decrease of 0.0719 in LPIPS, highlighting its superior performance across diverse real-world low-light conditions.
Qualitative Evaluation. Fig. 6 visually confirms the performance and drawbacks of the SLLIE methods on our LSD. Our model can generate visually plausible results with consistent brightness, better denoising, and less visible artifacts. Also, with and without noise, our TFFormer illustrates a consistent performance for enhancing complex low-light scenes.
@@ 4.2.1 Perceptual Comparison
Figure 7: Cross-dataset evaluation on LSD-U. Top row: low-light image enhanced by SLLIE dataset, bottom row: user rank of performance of the SLLIE method summarized by bredly-terry model.
The proposed LSD is designed to handle diverse real-world low-light scenarios effectively. We conducted the cross-dataset comparison to assess its practical applicability and user acceptance, evaluating enhancement results from both existing SLLIE datasets and our proposed LSD. Notably, LSD-U was collected independently of LSD to ensure an unbiased evaluation. Our cross-dataset evaluation aims to provide a fair and comprehensive comparison of SLLIE datasets rather than assessing performance on dataset-specific training distributions. Therefore, we trained our TFFormer model on SLLIE datasets and our LSD dataset. To evaluate model generalization on unknown, real-world scenes captured by various hardware, we conducted an extensive perceptual comparison and user study using our collected LSD-U benchmarking dataset. For this study, we invited 15 participants, each reviewing 5 out of 2,117 randomly selected image sets, ranking the images based on personal preference. We summarized these rankings using the Bradley-Terry model [1] to demonstrate model practicality for real-world SLLIE. As illustrated in Fig. 7, user preferences and perceptual comparisons indicate that our proposed method consistently outperformed the method trained on the existing SLLIE dataset.
Dataset/Class Bicycle Boat Bottle Bus Car Cat Chair Cup Dog Motorbike People Table mAP
@@
4.2.2 Vision Tasks
In addition to the visual enhancement, a robust SLLIE dataset can enhance vision tasks under stochastic lighting conditions. To assess the effectiveness of our LSD, we applied it alongside other SLLIE datasets to enhance the ExDark [28] validation dataset comprising 1,473 images and 13 object classes. We then evaluated object detection performance using the YOLOv3 model [10], trained on the COCO dataset. Tab. 3 shows that the proposed LSD improves detection accuracy by 6.80 mAP compared to the low-light ExDark inputs, outperforming other datasets.
@@ 4.4 Ablation Study
LSD. Fig. 9 shows TFFormer’s performance with filter patch training, full-scene training, and raw patch training. Our selective patch training approach allows TFFormer to learn local noise patterns better, preserving fine details in complex areas. In contrast, non-selective training often produces blurry outputs due to motion blur in the GT images, while full-scene training limits the model’s focus on intricate noise, emphasizing global attributes. Tab. 4 summarizes TFFormer’s performance on LSD (training) variants. Compared to the full-scene training, filter patches can improve TFFormer performance by 2.14 dB on PSNR.
Figure 9: Ablation study on LSD, patch and full-scene training
TFFormer In our ablation study, we evaluated each novel component of TFFormer by starting with a base model and progressively adding LCE, LCCAB, LCGRB, and LCPL. The LCE variant enhances denoising and color consistency, LCCAB improves feature correlation for global reconstruction, LCRGB refines the projected intermediate outputs, and LCP loss enhances the perceptual quality of the reconstructed image. TFFormer achieves the best results by combining all components, as shown in Tab. 5. Fig. 10 further confirms the contribution of the proposed component by visual comparison.
@@
LCE LCCAB LCGRB LCPL PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
Figure 10: Ablation study on the novel component of proposed TFFormer on LSD.
@@ 4.5 Discussion
The proposed method illustrates the practicability of a generic SLLIE method. Also, it reveals the limitations of existing SLLIE datasets and models. Please refer to the supplementary material for more details and results. Despite promising results, we observed that all deep models, including TFFormer, could struggle to produce visually appealing images in some extreme cases (e.g., NLI capture under 1 lux). Therefore, it has been planned that extreme low-light NLI enhancement will be studied separately through a future study. Also, joint OD with SLLIE could be an exciting choice for many industry applications.
@@ 5 Conclusion
This study proposed the largest SLLIE dataset collected with numerous hardware and sources. Our proposed LSD dataset contains 6,425 high-resolution low-light images with corresponding reference images. We collected these image pairs using dynamic calibration settings tailored to each device and capture condition. From these, we utilized 6,025 pairs to generate approximately 180,000 refined training patches and reserved an additional 400 full-scene images for SLLIE benchmarking. Furthermore, we collected 2,117 unique low-light scenes from diverse sources for aesthetic evaluations. Additionally, we proposed a novel deep model to learn SLLIE from collected data samples. Our proposed model leverages LC attributes in the transformer model to maintain color accuracy while denoising. We also incorporated an LCCAB, LCGRB, and LCPL to accelerate our SLLIE performance. We generalized our model for numerous hardware and extensively evaluated its practicability. It has been planned to learn joint SLLIE-OD and extreme NLI enhancement through a future study.
@@ C.2 Loss calculation
To perceive visually pluisable images, we leverage an additional L1 loss along with the proposed LCPL. Thereby, we defined final LC-guided loss as:
@@ D.1 TFFormer on LSD
D.2 LSD on Deep SLLIE Methods
@@
D.3 LSD-TFFormer on Real-world Scenes
@@
User Study. We also performed a separate user study to verify the mass acceptance of our LSD-TFFormer in real-world scenarios. In a user study, 35 evaluators aged 15 to 62 reviewed three random pairs per category (low-light vs. enhanced) based solely on aesthetic appeal, unaware of the study’s purpose. As shown in Tab. 13, over 82% of participants preferred images enhanced by LSD-TFFormer, with a particularly high preference for social media and older images, highlighting the model’s practical applicability to real-world SLLIE.
@@
D.4 Visual Improvement on Vision Task
@@
D.5 LSD-TFFormer vs Night mode

### v2

@@ 4.1 Comparison between SLLIE methods
Experimental Setup. First, we benchmarked SOTA methods on the proposed LSD. We included various SLLIE approaches: Retinex-based models (RetinexNet [42], Kind [54], Kind+[55], DeepUPE[40]), transformer-based methods (Retinexformer [3], IAT [6], SNRNet [45]), general image restoration models (Uformer [41], MIRNet [51]), a diffusion-based method (GSAD [13]), a hybrid diffusion-Retinex model (Diff-Retinex [50]), and non-RGB-based methods (LYT [2], HVI [47]). To ensure fair comparison, we reproduced and verified each method’s reported results on their recommended datasets, then retrained all models on the proposed LSD. For comprehensive evaluation, all models were trained on both DLI and NLI scenes.
Quantitative Evaluation. Tab. 2 shows the performance of SLLIE methods in different lighting and environments. Our proposed TFFormer achieves state-of-the-art results across all scenarios. Compared to the strong baseline RetinexFormer [3], TFFormer yields average gains of +2.04 dB in PSNR, +0.0738 in SSIM, and -0.0357 in LPIPS. It also surpasses the recent HVI [47] method by +2.26 dB in PSNR, +0.0859 in SSIM, and -0.0452 in LPIPS, confirming its robustness across diverse real-world lighting conditions.
In addition to accuracy, TFFormer offers a favorable performance complexity trade-off. With just 5.87M parameters and 34.6 GFLOPs, it is significantly more efficient than recent methods like Diff-Retinex [50] (80.7M, 492.5 GFLOPs) over 13$\times$ larger and 14$\times$ more computational yet achieves markedly lower performance. TFFormer outperforms it by +3.56 dB in PSNR, +0.2005 in SSIM, and -0.1206 in LPIPS. Compared to the heaviest model, MIRNet [51] (31.79M, 816 GFLOPs), TFFormer is 5.4$\times$ smaller and 24$\times$ faster, while achieving +3.83 dB higher PSNR, +0.1088 higher SSIM, and -0.0417 lower LPIPS. These results underline TFFormer’s efficiency without compromising quality
Qualitative Evaluation. Fig. 6 visually compares SLLIE methods on the LSD. Retinex-based models like Diff-Retinex [50] and Retinexformer [3] show color distortions. Diff-Retinex introduces yellow-green casts in NLI (Outdoor), while Retinexformer oversaturates warm tones in DLI (Outdoor). HVI [47] yields dim, desaturated outputs and fails to restore brightness and contrast, particularly in NLI (Indoor) environments. Structural fidelity also suffers: Diff-Retinex and HVI blur textures like wall gradients and bookshelf edges, and Retinexformer shows minor color bleeding that softens boundaries. In contrast, TFFormer recovers natural color, realistic brightness, and sharp structure, avoiding overenhancement in DLI and effectively denoising complex NLI scenes. The results confirm TFFormer’s robustness and generalization across both noisy and clean low-light scenes.
Figure 6: Qualitative comparison of existing SLLIE methods on the LSD dataset. The top scenes show performance on DLI, and the bottom example illustrates performance on DLI scenes. TFFormer produces cleaner, more plausible images, outperforming existing methods.
@@ 4.2.1 Perceptual Comparison
The proposed LSD is designed to address diverse real-world low-light scenarios. For fair assessment of generalization and user preference, we performed cross-dataset evaluation using the independently collected LSD-U set. Notably, LSD-U features a distribution entirely distinct from LSD and all existing SLLIE datasets. We trained TFFormer separately on existing SLLIE datasets and LSD, then evaluated their acceptance among mass users through a perceptual user study. We recruited 15 participants, each presented with 5 randomly selected sets from the 2,117-image LSD-U collection. Participants ranked the enhancement outputs based on visual preference. Rankings were aggregated using the Bradley-Terry model [1], revealing that TFFormer trained on LSD consistently outperformed counterparts trained on prior datasets (Fig. 7).
Figure 7: User rankings, summarized via the Bradley-Terry model, show a substantial margin preferred LSD outputs.
4.2.2 Vision Tasks
Beyond perceptual quality, a robust SLLIE dataset should facilitate downstream vision tasks in low-light conditions. To evaluate this, we enhanced the ExDark validation set [26], which includes 1,473 images across 13 object categories, using LSD and existing SLLIE datasets. We then performed object detection with YOLOv3 [9], pretrained on the COCO dataset [21]. As shown in Tab. 3, LSD-enhanced images achieved the highest average detection performance at 38.80 mAP, surpassing all other datasets. LSD also yielded the best or second-best results in most categories, notably improving detection for challenging classes such as Car, Cat, Chair, Cup, Dog, Motorbike, People, and Table. Compared to raw ExDark images, LSD-enhanced inputs provided a substantial +6.80 mAP gain, demonstrating the effectiveness in real-world low-light vision applications.
Dataset/Class Bicycle Boat Bottle Bus Car Cat Chair Cup Dog Motorbike People Table Avg.
@@ 4.4 Ablation Study
LSD. Fig. 9 and Tab. 4 show TFFormer’s performance with three training strategies: raw patches, full-scene images, and our proposed filtered patches. Filtered patch training yields the best results, improving PSNR by 2.14 dB on DLI and 1.62 dB on average compared to full-scene training. By focusing on structurally rich and clean regions, it enables the model to learn localized noise patterns and preserve fine details, avoiding spatial blurs introduced by motion in GT images in extreme cases. In contrast, raw patch training underperforms due to exposure to noisy or misaligned content, leading to degraded output quality.
Figure 9: Ablation study on LSD, patch, and full-scene training
@@
TFFormer. Tab. 5 and Fig. 10 present an ablation study evaluating each proposed component in TFFormer. Starting from a base model, we progressively add the decoupled LC Encoder (LCE), Cross-Attention (LCCAB), Refinement Block (LCGRB), and Consistency Guidance (LCCG). LCE enhances denoising and color consistency by separating luminance and chrominance early in the image processing pipeline. Adding LCCAB enables adaptive inter-domain fusion with joint decoding, boosting global reconstruction. LCGRB enhances structural detail recovery, while LCCG improves perceptual alignment. The full model achieves a +3.57 dB PSNR gain and a -0.0268 LPIPS reduction over the base, validating the effectiveness of our design for structure-aware and color-consistent enhancement in complex lighting conditions.
Figure 10: Ablation study on the novel component of the proposed TFFormer on LSD.
@@
LCE LCCAB LCGRB LCCG PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@ 4.5 Discussion
The proposed method validates a generic SLLIE framework and highlights limitations in existing datasets and models under diverse real-world conditions. Full implementation details, including sensor specs, app design, training strategies, and extended results, are provided in the supplementary material. While our model generalizes well, it struggles under extreme low-light (e.g., ¡1 lux), which we plan to address through joint denoising and enhancement in future work. We also aim to extend LSD to support downstream tasks like object detection and segmentation for practical applications.
@@ 5 Conclusion
We presented LSD, the largest real-world SLLIE dataset to date, comprising 6,425 high-resolution low-light image pairs captured using diverse devices and adaptive calibration strategies. From these, 6,025 pairs were used to generate approximately 180,000 training patches, with 400 full-resolution pairs reserved for benchmarking. Additionally, we collected 2,117 unpaired low-light images from varied sources to support aesthetic and generalization evaluations. To fully leverage this dataset, we proposed a novel Transformer-CNN model that decouples luminance and chrominance features for accurate color restoration and robust denoising. Our design integrates LCCAB, LCGRB, and LCPL modules to enhance global fusion, local refinement, and perceptual quality. Extensive evaluations demonstrate strong generalization across hardware, with future extensions aimed at joint SLLIE–object detection and extreme NLI enhancement.
@@ C.2 Loss calculation
To perceive visually plausible images, we leverage an additional L1 loss along with the proposed LCCG. Thereby, we defined final LC-guided loss as:
@@ D.1 TFFormer on LSD
D.2 TFFormer vs. Existing Method (More Results)
As we mentioned earlier, the challenges posed by over-enhancement, noise amplification, and color distortion in existing SLLIE methods. We present qualitative comparisons in Figures 21 and 22. These examples underscore TFFormer’s ability to generalize to complex, real-world scenes across varying illumination levels and conditions.
In Figure 21, we visually compare TFFormer against three representative SLLIE architectures: Diff-Retinex [50], HVI [47], and RetinexFormer [3], on both standard low-light (DLI) and challenging noisy low-light (NLI) scenarios. Despite leveraging Retinex priors [3, 50] or biologically inspired feature fusion [47], these baselines frequently suffer from overexposure, structural artifacts, and color distortions—especially under NLI conditions. In contrast, TFFormer consistently delivers results with more natural tone rendering and finer structure preservation, validating the effectiveness of its luminance–chrominance (LC) decoupling and guided refinement mechanism.
Figure 22 extends the comparison to a broader set of 10 state-of-the-art methods, including classical RetinexNet [42], Kind(+)[54, 55], DeepUPE[40], MIRNet [51], IAT [6], Uformer [41], SNRNet [45], GSAD [13], and LYT [2]. Across both DLI and NLI cases, TFFormer produces perceptually coherent, cleaner outputs with better structure and fewer color artifacts. Prior RGB-based transformer and CNN methods often suffer from over-smoothing (e.g., MIRNet, Uformer) or produce strong unnatural color casts (e.g., DeepUPE, SNRNet). Retinex-based designs, despite strong PSNR, tend to yield inconsistent tones under challenging lighting, likely due to excessive or misaligned enhancement. These observations, supported by both qualitative and quantitative evidence, underscore the robustness and generalization capacity of TFFormer in complex real-world scenes.
@@
Together, these visual results reinforce the paper’s central hypothesis: existing SLLIE models struggle to preserve structure and realism under challenging real-world scenarios, while TFFormer, trained on the proposed LSD dataset and guided by LC-aware modules, produces high-fidelity enhancements with superior generalization.
D.3 LSD on Deep SLLIE Methods
@@
D.4 LSD-TFFormer on Real-world Scenes
@@
User Study. We also performed a separate user study to verify the mass acceptance of our LSD-TFFormer in real-world scenarios. In this study, 35 participants aged 15 to 62 evaluated enhanced outputs across various lighting conditions and capture sources. Each participant was shown three randomly selected low-light versus enhanced image pairs from multiple SLLIE datasets and asked to choose the image they preferred, based solely on aesthetic appeal and without knowledge of the study’s purpose.
As summarized in Table 13, over 82% of participants preferred the outputs generated by LSD-TFFormer. The preference was particularly strong in categories involving older social media photos and noisy video frames, where traditional models tend to fail due to compression artifacts and low signal-to-noise ratios. This preference distribution supports the claim that our model generalizes beyond controlled lab settings and remains robust across highly diverse, real-world image sources.
Fig. 25 visually reinforces these findings by comparing LSD-trained models against those trained on existing datasets, such as LOL-V1/V2, LSRW, and NCLLIE. While models trained on traditional datasets often produce flat or over-smoothed results, LSD-trained TFFormer outputs exhibit more natural contrast, reduced artifacts, and better color fidelity. In many cases, the LSD-enhanced images preserved ambient lighting cues and semantic details that were lost in other reconstructions. This aligns with user preferences and demonstrates the practical utility of our dataset and model for consumer-grade image enhancement applications.
@@
D.5 Visual Improvement on Vision Task
@@
D.6 LSD-TFFormer vs Night mode